# Measure gradients, not activations! Enhancing neuronal activity in deep reinforcement learning

### v1

@@ 3.2 Gradient Magnitude based Neuronal Activity Metric (GraMa)
Learning activity is directly related to the gradient received by a neuron during training (Zhou and Ge, 2024). A natural and simple way to enhance ReDo-like metrics is to count how many neurons in each network layer fail to obtain meaningful gradients from the current sample batch. Given an input distribution $D$, let $|\nabla_{h_{i}^{l}}L(x)|$ denote the gradient magnitude of the neuron $i$ in layer $\ell$ under input $x\in D$ and $H^{\ell}$ be the number of neurons in layer $\ell$. Based on Equation 1, we compute a learning capacity score for each neuron using the normalized average of the corresponding layer $\ell$, as shown in Equation 2.
@@
GraMa determines that neuron $i$ in layer $\ell$ is inactive when $G_{i}^{\ell}\leq\tau$.
GraMa has the same form as the dormant neuron ratio, but redefines its core signal, using gradient magnitudes $|\nabla_{h_{i}^{l}}L(x)|$ rather than activation values $h_{i}^{l}(x)$. The pre-set threshold $\tau$ allows us to detect neurons with outlier gradient magnitude. Since gradient information is available in the tensor after backpropagation at each step, there is no need to store the intermediate outputs of each neuron during forward computation, which is required by activation-based metrics. This makes GraMa lightweight, as verified in Figure 5 (left).
@@
Neuron resetting is a simple technique widely used to preserve the learning capacity of deep RL agents (Nikishin et al., 2022). Our approach follows the ReDo pipeline  (Sokar et al., 2023), as outlined in Algorithm 1: during training, we periodically quantify the activity of neurons in all layers using GraMa; any neuron $i$ with $G_{i}^{\ell}\leq\tau$ is considered inactive and reinitialized. Specifically, reinitialization involves resetting incoming weights to the original weight distribution, while outgoing weights are set to zero.
@@
We control the number of reset neurons in ReGraMa to evaluate the performance of four agents with varying proportions of low learning capacity neurons on the Dog Walker task. Results in Figure 6 (a) indicate that performance degrades significantly as the ratio of low learning capacity neurons increases.
@@
We use GraMa with a threshold of $\tau=0.0095$ to sample $1000$ inactive neurons in the pre-training period of vanilla agent. We then trace the change of their learning capacity scores and analyze the score distribution over time, as shown in Figure 6 (b). The results indicate that none of the sampled neurons exceed the threshold of $0.0095$ as training progresses. This suggests that such neurons are unable to recover their learning ability independently, further underscoring the importance of resetting neurons with low learning capacity.
Figure 6: (a): Low learning capacity neurons have a direct negative impact on agent performance. X-axis represents the normalized performance, and the Y-axis denotes the BRO-nets with different proportions of low learning capacity neurons. (b): Neuron loss of learning capacity is irreversible. x-axis denotes the number of neurons under each GraMa score, and the y-axis shows the training phase. (define $>0.0095$ as active).
@@
We theoretically analyze the similarities between ReGraMa and ReDo in traditional architectures (MLP with ReLU activations (Sokar et al., 2023)) and draw the following conclusions:
Theorem 3.1.
@@
Proof. From the dormant neuron formula (Equation 1), we can conclude that:
@@
Since $|h_{i}^{\ell}(x)|\geq 0$, $|h_{i}^{\ell}(x)|=0\iff\;h_{i}^{\ell}(x)=0$. This means the neuron is almost never activated on the dataset $D$. The derivative of the ReLU (Agarap, 2018) activation function is:
@@
$\nabla_{h_{i}^{\ell}}f=\mathbb{E}_{x\in D}\left[\frac{\partial f}{\partial h_{i}^{\ell}(x)}\cdot\frac{\partial h_{i}^{\ell}(x)}{\partial h_{i}^{\ell}}\right]=\mathbb{E}_{x\in D}\left[\frac{\partial f}{\partial h_{i}^{\ell}(x)}\cdot\underbrace{\frac{\partial h_{i}^{\ell}(x)}{\partial z_{i}^{\ell}(x)}}_{=0}\cdot h^{\ell-1}(x)\right]=0.$
@@
We trained a traditional fully connected network with ReLU on the CIFAR-100 benchmark (Krizhevsky, 2009), following the continuous learning experimental setup in Dohare et al. (2024). Every 15 epochs, a new category of data is added to the training set, requiring the agent to classify samples from the whole data distribution. The two gray curves in Figure 5 (right) show that the vanilla agent’s accuracy gradually declines as training progresses, indicating a loss of learning ability.
@@
Meanwhile, the proportion of inactive neurons detected by GraMa increases over time and fluctuates with the same periodicity as the accuracy curve, with both stabilizing around epoch 150. To further validate ReGraMa, we conducted a two-part intervention study: (1) We reset neurons identified as inactive by the $\tau-$dormant neuron ratio (ReDo) whenever a new data category is introduced. The resulting pink curves in Figure 5 show that resetting dormant neurons can both improve performance and reduce GraMa ratio. (2) We then reset neurons flagged by ReGraMa as having low activity. This intervention produced improvements similar to those of ReDo, as illustrated by the orange curves in Figure 5. These results suggest that ReGraMa is as effective as the ReDo metric in identifying neuronal inactivity in simple network architectures.
When extending the context to complex architectures, resetting based on activation values can be less effective. We evaluated the effectiveness of both metrics by pruning those neurons from a complex BRO-net agent and analyzing the impact on performance. The degree of fluctuation in accuracy reflects the importance of the pruned neurons. As shown in Figure 7, pruning neurons identified by ReGraMa has minimal effect on performance, confirming that gradient magnitude is a robust measure of neuronal inactivity in complex architectures.
Figure 8: (Top row) performance of each reset method with default BRO-net size. (Bottom row) inactive neuron population. Averages over four seeds.
@@ 4 Experiments
We conduct a series of experiments to investigate whether ReGraMa can mitigate neuronal activity loss and enhance performance. Specifically, we evaluate the effectiveness of ReGraMa across three representative and widely adopted architecture types: (i) the residual network-based policy (subsection 4.1), (ii) the online policy parameterized by a diffusion model (subsection 4.2), and (iii) the MLP policy featuring various activation functions (subsection 4.3). Finally, we verify the robustness of ReGraMa with respect to the threshold $\tau$ (Appendix B).
@@ 4.1 Residual Net-based Policy
Recent studies (Nauman et al., 2024b, Lee et al., 2025a) have shown that integrating residual modules into deep RL agents can significantly improve representation capability on complex visual tasks. We choose BRO-net (Nauman et al., 2024b) as a representative baseline and evaluate all methods on four challenging tasks from the DeepMind Control Suite (Tassa et al., 2018). All the algorithm parameters follow the default settings. We set the empirical threshold $\tau=0.01$ for ReGraMa, and use the same ReDo’s hyperparameters. The reset period is fixed at $1000$ steps. Further details are provided in Appendix A.1.
Figure 9: ReGraMa is robust under network scaling. Results averaged over 4 DMC tasks with 12 seeds per method.
@@
Results in Figure 8 show that ReGraMa can accurately reset the neurons with low activity in each stream of the multi-branch network, thus effectively maintaining the learning capacity of the deep RL agent on the four complex tasks and enabling continual learning. In contrast, ReDo performs poorly. This performance gap arises because, as discussed in subsection 3.1, ReDo misidentifies inactive neurons in multi-branch networks, undermining the effectiveness of its reset schedule.
@@
To assess scalability, we evaluate both ReDo and ReGraMa under increased network depth using BRO-net across four challenging environments: Dog Walk, Dog Stand, Humanoid Walk, and Humanoid Stand. As shown in Figure 9, ReGraMa maintains a stable advantage even as network size grows. In contrast, ReDo exhibits limited improvement and even degrades performance in the 2$\times$ model.
@@ 4.2 Diffusion Model-Based Policy
Figure 11: Proportion of inactive neurons in two MuJoCo tasks.
Figure 12: (Top row) performance under different activation functions. (Bottom row) inactive neuron population. Curves are averaged over four seeds.
Recent works have demonstrated that diffusion models, due to their strong expressiveness over multi-modal distributions, can significantly improve RL performance on complex control tasks (Chi et al., 2023). We use DACER (Wang et al., 2024), a recent online diffusion policy, as a baseline and test ReDo and ReGraMa on two MuJoCo-v4 tasks from the original paper. DACER relies on a Unet backbone with Swish activations. All parameters follow official defaults. Hyperparameters for ReGraMa and ReDo are the same as in subsection 4.1, and reset period is fixed at 1,000 steps. Full details are in Appendix A.2.
@@ Results.
As shown in Figure 11, ReGraMa maintains robust and consistent performance on this complex architecture, whereas ReDo provides only marginal improvements. This discrepancy is explained by the ratio of inactive neurons in Figure 11, which reveals that ReDo fails to reset neurons that have lost learning capacity. In contrast, ReGraMa accurately identifies and resets low-activity neurons, preserving the agent’s ability to learn and improving overall performance.
@@ 4.3 Activation Function Variants
Nadimpalli et al. (2025) shows that saturated activation functions in the hidden layers may improve RL agents’ performance. Following their setup, we replace ReLU in SAC with Tanh and Sigmoid, keeping all other parameters as default. This enables us to evaluate the robustness of ReGraMa across various activation functions while minimizing the influence of extraneous factors on the experimental outcomes. We set $\tau=0$ for ReGraMa, and configure ReDo as in subsection 4.1. All methods are evaluated on the challenging Ant task. Hyperparameter are provided in Appendix A.3.
@@ Results.
Results in Figure 12 show that ReGraMa accurately identifies low-quality neurons, mitigating neuronal inactivity and avoiding instability caused by false reset. While ReDo performs well under the ReLU, its performance degrades significantly with other activation functions, sometimes failing below vanilla SAC. In Figure 12(c), although ReGraMa outperforms ReDo under the Tanh activation function, the proportion of inactive neurons gradually increases during training. We leave further investigation into whether this behavior stems from the perspective of the activation function or reset mechanism to future work.
5 Related Work
Dynamic objectives may cause irreversible damage to the neuronal activity of networks during training (Lyle et al., 2023, Nauman et al., 2024a), which is also significant in the field of multi-agent RL (Qin et al., 2024) and visual deep RL (Ma et al., 2024). This issue may lead to the phenomenon where deep learning agents progressively lose their ability to fit new data (Abbas et al., 2023), and limit their capacity for continual learning (Elsayed and Mahmood, 2024). In the field of Deep RL, some factors have been found to have a direct correlation with the activity of neurons. For instance, substituting the ReLU activation has been shown to preserve neuron-level learning capacity (Abbas et al., 2023). Certain loss functions, especially those incorporating L2 regularization, have been found to mitigate the network’s activity degradation (Kumar et al., 2023). Another relevant line of research explores architectural adjustments such as scaling up (Nikishin et al., 2023) or topology growth (Liu et al., 2025, Evci et al., 2020, Tan et al., 2022). These methods can introduce modules that enhance the proportion of active, high-quality neurons.
A separate and increasingly influential line of research focuses on directly manipulating model parameters to recover their activity. Resetting neural networks, either periodically or selectively, has been shown to be both effective and easy to implement (Farias and Jozefiak, 2025, Yun et al., 2025). For instance, Nikishin et al. (2022) demonstrates the benefits of periodic resets for continual learning. Similarly, Ma et al. (2024) shows that selectively resetting specific layers can improve stability without sacrificing expressive capacity. At a finer granularity, ReDO (Sokar et al., 2023) introduces a neuron-level resetting scheme that outperforms earlier coarse strategies in terms of stability and precision. However, ReDO’s reliance on activation-based quantization limits its applicability to more complex architectures (Lyle et al., 2023, 2024). To address this limitation, we introduce a novel neuronal activity quantification approach based on gradient magnitudes, enabling generalized activity estimation and recovery across arbitrary network designs.
@@ 6 Discussion
This research focuses on the critical issue of neuronal activity loss during training in deep RL agents. We show that the commonly used $\tau$-dormant neuron ratio (ReDo), which relies on neuron activations, struggles to capture learning activity in modern, highly parameterized agents. This limitation arises because activation sparsity does not directly reflect a neuron’s contribution to learning. We shift focus from activations to gradients, and introduce GraMa, a simple, efficient, and architecture-agnostic metric based on gradient magnitude. GraMa enables accurate tracking of neuron-level learning dynamics across a broad range of network types, including residual and diffusion-based policies, and substantially recovers learning activity via guiding neuron resetting (subsection 3.2). Our findings suggest that even high-capacity policies in deep RL suffer from underutilization at the neuron level, which may hinder generalization, multi-task transfer, and continual adaptation as in supervised learning (Dohare et al., 2024). By providing a lightweight diagnostic and intervention tool, GraMa opens the door to more principled approaches for maintaining continuous learning ability in deep RL agents.
@@
Our evaluation is limited to three representative neural architectures due to computational constraints. Future work will extend GraMa to more complex settings, including large-scale transformer policies and multi-task environments. We aim for GraMa to serve as a practical tool for diagnosing and preserving learning capacity in deep RL agents.
7 Acknowledgment
This work is supported by the National Natural Science Foundation of China 62406266. We would also like to thank the Python community (Van Rossum and Drake Jr, 1995, Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020) and Matplotlib (Hunter, 2007).

### v2

@@ 4.2 Gradient Magnitude based Neuronal Activity Metric (GraMa)
Learning activity is directly related to the gradient received by a neuron during training (zhou2024how). A natural and simple way to enhance ReDo-like metrics is to count how many neurons in each network layer fail to obtain meaningful gradients from the current sample batch. Given an input distribution $D$, let $|\nabla_{h_{i}^{l}}L(x)|$ denote the gradient magnitude of the neuron $i$ in layer $\ell$ under input $x\in D$ and $H^{\ell}$ be the number of neurons in layer $\ell$. Based on Eq. 1, we compute a learning capacity score for each neuron using the normalized average of the corresponding layer $\ell$, as shown in Eq. 2.
@@
GraMa determines that neuron $i$ in layer $\ell$ is inactive when $G_{i}^{\ell}\leq\tau$. GraMa has the same form as the dormant neuron ratio, but redefines its core signal, using gradient magnitudes $|\nabla_{h_{i}^{l}}L(x)|$ rather than activation values $h_{i}^{l}(x)$. The pre-set threshold $\tau$ allows us to detect neurons with outlier gradient magnitude. Since gradient information is available in the tensor after backpropagation at each step, there is no need to store the intermediate outputs of each neuron during forward computation, which is required by activation-based metrics. This makes GraMa lightweight, as verified in Fig. 5 (left).
@@
Neuron resetting is a simple technique widely used to preserve the learning capacity of deep RL agents (Nikishin2022ThePB). Our approach follows the ReDo pipeline  (sokar2023dormant), as outlined in Algorithm 1: during training, we periodically quantify the activity of neurons in all layers using GraMa; any neuron $i$ with $G_{i}^{\ell}\leq\tau$ is considered inactive and reinitialized. Specifically, reinitialization involves resetting incoming weights to the original weight distribution, while outgoing weights are set to zero. It is worth noting that, empirically, we found that although LayerNorm layers contain only two learnable parameters, they tend to be quite inactive. As a result, ReGraMa is deployed for these parameters in our official implementation, leading to performance improvements that warrant further analysis. Another avenue worth exploring is the potential for resetting the input weights and biases of neurons.
@@
We control the number of reset neurons in ReGraMa to evaluate the performance of four agents with varying proportions of low learning capacity neurons on the Dog Walker task. Results in Fig. 6 (a) indicate that performance degrades significantly as the ratio of low learning capacity neurons increases.
Figure 6: (a): Low learning capacity neurons have a direct negative impact on agent performance. X-axis represents the normalized performance, and the Y-axis denotes the BRO-nets with different proportions of low learning capacity neurons. (b): Neuron loss of learning capacity is irreversible. x-axis denotes the number of neurons under each GraMa score, and the y-axis shows the training phase. (define >0.0095>0.0095 as active).
@@
We use GraMa with a threshold of $\tau=0.0095$ to sample $1000$ inactive neurons in the pre-training period of vanilla agent. We then trace the change of their learning capacity scores and analyze the score distribution over time, as shown in Fig. 6 (b). The results indicate that none of the sampled neurons exceed the threshold of $0.0095$ as training progresses. This suggests that such neurons are unable to recover their learning ability independently, further underscoring the importance of resetting neurons with low learning capacity.
@@
We theoretically analyze the similarities between ReGraMa and ReDo in traditional architectures (MLP with ReLU activations (sokar2023dormant)) and draw the following conclusions:
Theorem 1.
@@
Proof. From the dormant neuron formula (Eq. 1), we can conclude that:
@@
Since $|h_{i}^{\ell}(x)|\geq 0$, $|h_{i}^{\ell}(x)|=0\iff\;h_{i}^{\ell}(x)=0$. This means the neuron is almost never activated on the dataset $D$. The derivative of the ReLU (Agarap2018DeepLU) activation function is:
@@
$\nabla_{z_{i}^{\ell}}f=\mathbb{E}_{x\in D}\left[\frac{\partial f}{\partial h_{i}^{\ell}(x)}\cdot\frac{\partial h_{i}^{\ell}(x)}{\partial h_{i}^{\ell}}\right]=\mathbb{E}_{x\in D}\left[\frac{\partial f}{\partial h_{i}^{\ell}(x)}\cdot\underbrace{\frac{\partial h_{i}^{\ell}(x)}{\partial z_{i}^{\ell}(x)}}_{=0}\cdot h^{\ell-1}(x)\right]=0.$
@@
We trained a traditional fully connected network with ReLU on the CIFAR-100 benchmark (Krizhevsky2009LearningML), following the continuous learning experimental setup in Dohare2024LossOP. Every 15 epochs, a new category of data is added to the training set, requiring the agent to classify samples from the whole data distribution. The two gray curves in Fig. 5 (right) show that the vanilla agent’s accuracy gradually declines as training progresses, indicating a loss of learning ability. Meanwhile, the proportion of inactive neurons detected by GraMa increases over time and fluctuates with the same periodicity as the accuracy curve, with both stabilizing around epoch 150. To further validate ReGraMa, we conducted a two-part intervention study: (1) We reset neurons identified as inactive by the $\tau$-dormant neuron ratio (ReDo) whenever a new data category is introduced. The resulting pink curves in Fig. 5 show that resetting dormant neurons can both improve performance and reduce GraMa ratio. (2) We then reset neurons flagged by ReGraMa as having low activity. This intervention produced improvements similar to those of ReDo, as illustrated by the orange curves in Fig. 5. These results suggest that ReGraMa is as effective as the ReDo metric in identifying neuronal inactivity in simple network architectures.
@@
When moving to more complex architectures, resetting neurons based solely on activation values becomes less reliable. In deep networks with normalization layers, residual connections, or context-dependent features, activations can fluctuate substantially even for neurons that are not meaningfully contributing to learning. As a result, activation magnitude alone is a noisy indicator of long-term inactivity. To examine this limitation, we compared activation-based and gradient-based criteria by pruning the corresponding neurons from a complex BRO-net agent and evaluating the resulting performance.
The degree of post-pruning performance degradation reflects both the relevance of the pruned neurons and the reliability of the underlying metric. As shown in Fig. 7, pruning neurons identified by ReGraMa causes only minimal performance loss, indicating that gradient magnitude provides a more stable and task-relevant measure of neuronal inactivity. This finding suggests that ReGraMa detects structural inactivity—neurons that remain consistently uninformative during optimization—rather than transient activation noise. Such stability is particularly advantageous in large, modular architectures, where distinguishing genuine inactivity from context-specific silence is critical for effective pruning and interpretability.
@@ 5 Experiments
We conduct a series of experiments to investigate whether ReGraMa can mitigate neuronal activity loss and enhance performance. Specifically, we evaluate the effectiveness of ReGraMa across three representative and widely adopted architecture types: (i) the residual network-based policy (Sec. 5.1), (ii) the online policy parameterized by a diffusion model (Sec. 5.2), and (iii) the MLP policy featuring various activation functions (Sec. 5.3). Finally, we verify the robustness of ReGraMa with respect to the threshold $\tau$ (Appendix B).
Figure 8: Performance and neuron inactivity with the default BRO-net size. (Top row) Episode return across four environments (Humanoid Stand, Dog Walk, Dog Stand, Humanoid Walk). (Bottom row) Corresponding proportion of inactive neurons. ReGraMa consistently achieves higher returns while maintaining fewer inactive neurons compared to ReDo and the vanilla baseline, demonstrating its effectiveness in stabilizing learning dynamics. Results are averaged over four seeds.
@@ 5.1 Residual Net-based Policy
Recent studies (nauman2024bigger; lee2025simba) have shown that integrating residual modules into deep RL agents can significantly improve representation capability on complex visual tasks. We choose BRO-net (nauman2024bigger) as a representative baseline and evaluate all methods on four challenging tasks from the DeepMind Control Suite (Tassa2018DeepMindCS). All the algorithm parameters follow the default settings. We set the empirical threshold $\tau=0.01$ for ReGraMa, and use the same ReDo’s hyperparameters. The reset period is fixed at $1000$ steps. Further details are provided in Appendix A.1.
@@
Results in Fig. 8 show that ReGraMa can accurately reset the neurons with low activity in each stream of the multi-branch network, thus effectively maintaining the learning capacity of the deep RL agent on the four complex tasks and enabling continual learning. In contrast, ReDo performs poorly. This performance gap arises because, as discussed in Sec. 4.1, ReDo misidentifies inactive neurons in multi-branch networks, undermining the effectiveness of its reset schedule.
Figure 9: ReGraMa is more robust under network scaling. Results averaged over four DMC tasks with 12 seeds per method.
@@
To assess scalability, we evaluate both ReDo and ReGraMa under increased network depth using the BRO-net architecture across four challenging environments: Dog Walk, Dog Stand, Humanoid Walk, and Humanoid Stand. As the network grows deeper, training stability and gradient signal propagation typically become harder to maintain, often leading to degraded performance or inefficient utilization of added capacity. In this context, we analyze whether the gradient-based reset criterion of ReGraMa remains effective under larger model scales.
[…]
@@ 5.2 Diffusion Model-Based Policy
Recent works have demonstrated that diffusion models, due to their strong expressiveness over multi-modal distributions, can significantly improve RL performance on complex control tasks (Chi2023DiffusionPV). We use DACER (wang2024diffusion), a recent online diffusion policy, as a baseline and test ReDo and ReGraMa on two MuJoCo-v4 tasks from the original paper. DACER relies on a Unet backbone with Swish activations. All parameters follow official defaults. Hyperparameters for ReGraMa and ReDo are the same as in Sec. 5.1, and the reset period is fixed at 1,000 steps. Full details are in Appendix A.2.
Figure 11: Proportion of inactive neurons during training across two MuJoCo tasks (Ant and Walker2d). ReGraMa maintains a consistently lower ratio of inactive neurons.
@@ Results.
As shown in Fig. 10, ReGraMa maintains robust and consistent performance on this complex architecture, whereas ReDo provides only marginal improvements. This discrepancy is explained by the ratio of inactive neurons in Fig. 11, which reveals that ReDo fails to reset neurons that have lost learning capacity. In contrast, ReGraMa accurately identifies and resets low-activity neurons, preserving the agent’s ability to learn and improving overall performance.
@@ 5.3 Activation Function Variants
nadimpalli2025evolving shows that saturated activation functions in the hidden layers may improve RL agents’ performance. Following their setup, we replace ReLU in SAC with Tanh and Sigmoid, keeping all other parameters as default. This enables us to evaluate the robustness of ReGraMa across various activation functions while minimizing the influence of extraneous factors on the experimental outcomes. We set $\tau=0$ for ReGraMa, and configure ReDo as in Sec. 5.1. All methods are evaluated on the challenging Ant task. Hyperparameter are provided in Appendix A.3.
@@ Results.
Results in Fig. 12 show that ReGraMa accurately identifies low-quality neurons, mitigating neuronal inactivity and avoiding instability caused by false reset. While ReDo performs well under the ReLU, its performance degrades significantly with other activation functions, sometimes failing below vanilla SAC. In Fig. 12(c), although ReGraMa outperforms ReDo under the Tanh activation function, the proportion of inactive neurons gradually increases during training. We leave further investigation into whether this behavior stems from the perspective of the activation function or reset mechanism to future work.
Figure 12: Effect of activation functions on performance and neuron inactivity. (Top row) Episode return across training under different activation functions (ReLU, Sigmoid, Tanh). (Bottom row) Corresponding proportion of inactive neurons. ReGraMa maintains higher performance and lower inactivity levels across all activations, indicating that gradient-based resetting generalizes effectively beyond specific nonlinearities. Curves are averaged over four seeds.
@@ 6 Discussion
This research focuses on the critical issue of neuronal activity loss during training in deep RL agents. We show that the commonly used $\tau$-dormant neuron ratio (ReDo), which relies on neuron activations, struggles to capture learning activity in modern, highly parameterized agents. This limitation arises because activation sparsity does not directly reflect a neuron’s contribution to learning. We shift focus from activations to gradients, and introduce GraMa, a simple, efficient, and architecture-agnostic metric based on gradient magnitude. GraMa enables accurate tracking of neuron-level learning dynamics across a broad range of network types, including residual and diffusion-based policies, and substantially recovers learning activity via guiding neuron resetting (Sec. 4.2). Our findings suggest that even high-capacity policies in deep RL suffer from underutilization at the neuron level, which may hinder generalization, multi-task transfer, and continual adaptation as in supervised learning (Dohare2024LossOP). By providing a lightweight diagnostic and intervention tool, GraMa opens the door to more principled approaches for maintaining continuous learning ability in deep RL agents.
@@
Our evaluation is limited to three representative neural architectures due to computational constraints. Future work will extend GraMa to more complex settings, including large-scale transformer policies, multi-task (mclean2025mu; nauman2025bigger), and continual deep RL (tang2025mitigating). We aim for GraMa to serve as a practical tool for diagnosing and preserving learning capacity in deep RL agents.
Acknowledgment
This work is supported by the National Natural Science Foundation of China 62406266. The authors would like to thank the reviewers for providing valuable feedback on the paper. We would also like to thank the Python community van1995python; 4160250 for developing tools that enabled this work, including NumPy harris2020array, Matplotlib hunter2007matplotlib, Jupyter 2016ppap, and Pandas McKinney2013Python.
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.